The probability that out of 10 persons, all born in April, at least two have the same birthday is:
Answer & explanation
Correct answer: option 4
Total no. of ways to allot 30 days to 10 people = 3010
Allot 10 days to 10 people as birthday = 30C10 × 10!
P(at least 2 have same birthday) = 1 - P (no one has same birthday) = $1-\frac{{^{30}C}_{10}×10!}{30^{10}}$